in a triangle abc angle b is 4 times angle a and angle c is 3 degrees less than 5 times angle A
Answer by stanbon(75887)

in a triangle abc angle b is 4 times angle a and angle c is 3 degrees less than 5 times angle A
------
Let size of angle a be "x"
The b = 4x
And c = 5x-3
------------------
Equation:
x + 4x + 5x-3 = 180 degrees
10x = 183
x = 18.3 degrees
----
a = 18.3 degrees
b = 4(18.3) degrees
c = 5(18.3)-3 degrees
========================
Cheers,
Stan H.